# Adjusting for Heavy Censoring and Double-Dipping to Compare Risk Stratification Abilities of Existing Models for Time to Diagnosis of Huntington Disease

### v1

@@ Abstract
Huntington disease (HD) is a genetically inherited neurodegenerative disease with progressively worsening symptoms. Accurately modeling time to HD diagnosis is essential for clinical trial design and treatment planning. Langbehn’s model, the CAG-Age Product (CAP) model, the Prognostic Index Normed (PIN) model, and the Multivariate Risk Score (MRS) model have all been proposed for this task. However, differing in methodology, assumptions, and accuracy, these models may yield conflicting predictions. Few studies have systematically compared these models’ performance, and those that have could be misleading due to (i) testing the models on the same data used to train them and (ii) failing to account for high rates of right censoring ($80\%+$) in performance metrics. We discuss the theoretical foundations of the four most common models of time to HD diagnosis, offering intuitive comparisons about their practical feasibility. Further, we externally validate their risk stratification abilities using data from the ENROLL-HD study and performance metrics that adjust for censoring. Our findings guide the selection of a model for HD clinical trial design. The MRS model, which incorporates the most covariates, performed the best. However, the simpler CAP and PIN models were not far behind and may be logistically simpler to adopt. We also show how these models can be used to estimate sample sizes for an HD clinical trial, emphasizing that previous estimates would lead to underpowered trials.
Keywords Brier score $\cdot$ Censored covariates $\cdot$ Cox model $\cdot$ Inverse probability of censoring weights $\cdot$ Neurodegenerative disease $\cdot$ ROC curves
@@ 1.1 Huntington Disease, Genetic Testing, and Observational Studies
Huntington disease (HD) is a hereditary neurodegenerative disease caused by a repeat expansion of the trinucleotide CAG (cytosine-adenine-guanine) in the Huntingtin gene MacDonald1993. This disease is fully penetrant for patients with $\geq 40$ CAG repeats, meaning that they will meet the diagnostic criteria for Huntington disease with probability $90\%$ Tabrizi2022b. An HD diagnosis is made based on a clinician’s assessment of a patient’s motor impairment, following the Unified Huntington Disease Rating Scale (UHDRS) Kremer1996.
The full penetrance of HD is unique among neurodegenerative diseases. With genetic testing, researchers can identify patients who will develop HD before symptom onset and construct valuable observational cohorts (see Table 2 for examples). These studies all collected crucial clinical measurements of HD symptom severity. The UHDRS Diagnostic Confidence Level (DCL) is a categorical rating from 0–4 assigned by a clinician, where a DCL of 0 indicates no motor impairment and 4 indicates that the clinician is 99% certain that the patient is exhibiting motor impairment due to HD Paulsen2019. Crucially, a DCL rating of 4 indicates a clinical diagnosis. The UHDRS Total Motor Score (TMS) Kremer1996 is a 31-item, clinician-rated assessment of motor impairment, where higher scores indicate greater impairment. The Symbol Digit Modalities Test (SDMT) Smith1973 and Stroop Word, Color, and Interference tests Stroop1935 are timed assessments that quantify cognitive ability, with lower scores indicating worse cognitive impairment.
Table 1: Summary of major Huntington disease observational studies.
Study Years Sample Size Right-Censoring Rate Citation
COHORT 2005–2011 $1326$ 81% Dorsey and Investigators (2012) Dorsey2012
TRACK-HD 2008–2011 257 85% Tabrizi et al. (2009) Tabrizi2009
PREDICT-HD 2002–2014 $1159$ 77% Paulsen et al. (2006) Paulsen2006
ENROLL-HD 2012–present $5173$ 88% Sathe et al. (2021) Sathe2021
Table 2: Summary of Huntington disease studies, including study name, years active, sample size, and right-censoring rate (the percent of patients undiagnosed by the end of study).
@@ 1.2 Designing Clinical Trials: Traditional Versus Preventative
Clinically, HD is characterized by progressive motor dysfunction, cognitive decline, and loss of functional independence. Numerous pharmacological, psychiatric, and behavioral treatments have been proposed for HD patients, but none so far have been approved for delaying or stopping disease progression. However, several studies are in progress and one has shown promise in its human trials to reduce Huntington disease progression UCL2025_AMT130. Another focus is on palliative care and improving quality of life, but treatments remain limited. Chorea, characterized by involuntary movements, is the only HD symptom with an approved pharmacological treatment OlmedoSaura2025. Approved symptom treatments and recent clinical breakthroughs further highlight the need for early intervention. Risk stratification models are key to designing preventative clinical trials and identifying patients before symptoms manifest, allowing for potential interventions to have the most effect before irreversible symptoms and damage are incurred.
@@
Traditionally, clinical trials for HD recruit patients who have already received their clinical diagnosis, and are thus displaying severe motor, cognitive, and functional impairment. This recruitment strategy allows for shorter trials and increases the likelihood of detecting treatment effects, which are both desirable traits to trial coordinators Langbehn2020. However, this approach has a major limitation: By the time of diagnosis (and thus trial recruitment), neurological damage is irreversible and therapeutic interventions may be ineffective. Moreover, neurodegenerative diseases like HD progress slowly for decades prior to clinical diagnosis. Thus, trials that only recruit subjects after their clinical diagnosis could fail to detect interventions that impact the disease’s early stages.
@@
Given these limitations, researchers have begun to shift clinical trial focus to patients with prodromal HD, that is, those who have not yet been diagnosed but are genetically guaranteed to develop the disease Rubinsztein2016; Langbehn2020. These studies, called preventative clinical trials, aim to slow or stop symptom development and thus delay or prevent the clinical diagnosis of a patient with HD.
@@
However, one unique challenge with preventative trials is the small proportion of patients expected to be diagnosed during the trial; Paulsen et al. (2019) Paulsen2019 report that less than $1\%$ of patients in an observational study of prodromal HD patients were diagnosed within 3 years, a typical trial length. Since statistical power is tied to the number of “events” (HD diagnoses) in the untreated group, these trials often require large sample sizes to be sufficiently powered to detect a treatment effect. Put simply, when fewer patients are expected to be diagnosed during a trial, more must be enrolled to detect a treatment effect. With larger sample sizes, these trials can become more expensive and difficult to implement.
@@
Fortunately, time-to-diagnosis models can help to overcome this challenge via a strategy known as sample enrichment Zhang2011; Long2017; Paulsen2019; Langbehn2022. We can use them to calculate log-risk scores for potential patients and recruit only those whose log-risk scores lie above a particular threshold (i.e., who are closer to diagnosis). This approach increases the expected rate of diagnosis in the untreated group, reducing the sample size needed for a desired statistical power. In doing so, sample enrichment makes preventative clinical trials for HD more cost-efficient and easier to implement.
@@ 1.3 Comparing Existing Models for Time to Diagnosis
Time-to-diagnosis models for HD are essential tools for both clinical trial planning and patient counseling. These models help estimate when an at-risk patient (based on genetic testing) will receive a clinical diagnosis. Currently, there exist four major time-to-diagnosis models for HD: Langbehn Langbehn2004, CAG-Age Product (CAP) Zhang2011, Multivariate Risk Score (MRS) Long2015, and Prognostic Index Normed (PIN) Long2017. All four are survival models built using data from observational studies of prodromal patients ($\geq 36$ CAG repeats), as described in Section 2.2.
@@
As these models were developed, some comparisons were made to the existing models, but none considered all four. The Langbehn model was developed first, so there were no comparisons to make Langbehn2004. The CAP model came next, and Zhang et al. (2011) Zhang2011 discussed the Langbehn model but did not provide quantitative comparisons. Long and Paulsen (2015) Long2015 compared the newly-proposed MRS model to a nonparametric model similar to CAP but again made no comparison to Langbehn. Finally, Long et al. (2017) Long2017 compared the PIN model to a semiparametric model similar to CAP, but made no comparison to Langbehn or MRS.
Despite the public availability of these modeling tools, limited work has evaluated and compared them systematically. One of the few comparisons comes from Paulsen et al. (2019) Paulsen2019, who assessed the predictive performance of the CAP, MRS, and PIN models using receiver operating characteristic (ROC) analysis. In short, this technique evaluates how well a model separates patients who will be diagnosed up to versus after a specific time. There are three key limitations to the model evaluation conducted by Paulsen et al. (2019) Paulsen2019.
First, only three of the four major models were assessed. Second, their ROC analyses did not account for the high proportion of patients in their data who remained undiagnosed at study end (i.e., whose times to diagnosis were right censored). Although the time-to-diagnosis models were designed to handle censored data, this validation approach was not. Standard ROC analyses assume that all patients’ diagnosis times are known. Third, there was double dipping Ball2020, because the models were evaluated on the PREDICT-HD dataset Paulsen2006 from which they were built. (The CAP and MRS models were developed from PREDICT-HD alone; the PIN model was developed on a combination of studies that included PREDICT-HD.)
Additional work by Long et al. (2023) examined the PIN and CAP models in the context of the Huntington Disease Integrated Staging System (HD-ISS) Long2023, which can be used as an alternative outcome measure to time to clinical diagnosis in Huntington disease clinical trials Tabrizi2022b. Long et al. (2023) Long2023 also used ROC analysis to compare model performance, incurring similar limitations to Paulsen et al. (2019) Paulsen2019 due to censoring. Moreover, their comparisons excluded the Langbehn and MRS models, and performance in predicting HD-ISS does not generalize to other HD outcomes.
@@ 2.1 Notation
Suppose we observe data for $n$ patients, indexed by i=1,,ni=1,...,n. Let $X_{i}$ denote patient $i$’s continuous diagnosis time and $\boldsymbol{Z}_{i}$ denote their covariate vector. In survival analysis, our goal is typically to estimate some aspect of the conditional distribution of $X_{i}$ given $\boldsymbol{Z}_{i}$. When the primary goal is sample enrichment, the log-risk score for $X_{i}$ given $\boldsymbol{Z}_{i}$ is the estimand of interest, denoted by ${\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}$ for some model-specific parameter vector $\boldsymbol{\beta}$} that describes this conditional distribution. This score is a scalar summary of patient $i$’s risk of being diagnosed with HD, where higher scores indicate greater risk Park2021.
However, analysts often encounter data in which $X_{i}$ is censored by another variable $C_{i}$ for some patients. We focus on random right censoring, where $X_{i}$ is censored by the random variable $C_{i}$ if $X_{i}>C_{i}$. In many HD studies, more than 75% of patients’ diagnosis times are randomly right censored. Typically, censored patients dropped out of a study before being diagnosed or the study ended before they were diagnosed, in which case $C_{i}$ would be time to drop out or study end, respectively. Hence, we have $W_{i}=\min(X_{i},C_{i})$ and $\Delta_{i}=\textrm{I}(X_{i}\leq C_{i})$ so that our observed, patient-level data are $(W_{i},\Delta_{i},\boldsymbol{Z}_{i})$. Our goal, then, is to estimate log-risk scores ${\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}$ using these data, though our true interest is in the full data pair $(X_{i},\boldsymbol{Z}_{i})$.
@@ 2.2.1 Model 1: Langbehn
The Langbehn model was introduced to estimate the probability of HD diagnosis at a given age as a function of CAG repeats Langbehn2004. It was developed using retrospective data collected from 40 clinical sites across Africa, Asia, Europe, and North America. While this dataset did not come from one of the commonly cited cohorts in Table 2, it constituted the largest dataset of HD patients to date.
When developing this model, Langbehn et al. (2004) Langbehn2004 restricted the dataset to patients with $41$ to $56$ CAG repeats (inclusive). After these exclusions, they had $2913$ patients, of whom $2298$ had been clinically diagnosed by study end. With 615 patients who had not been diagnosed (i.e., whose times of diagnosis had been right-censored), the sample saw a right-censoring rate of $21\%$. This proportion contrasts sharply with the cohorts described in Table 2, which saw between $77\%$ and $88\%$ undiagnosed patients.
@@
where $\mu(\textrm{CAG}_{i};{\boldsymbol{\beta}})=\beta_{0}+\exp(\beta_{1}-\beta_{2}\textrm{CAG}_{i})$ is the mean age at diagnosis and $\sigma^{2}(\textrm{CAG}_{i};{\boldsymbol{\gamma}})={\gamma_{0}+\exp(\gamma_{1}-\gamma_{2}\textrm{CAG}_{i})}$ its variance (both dependent on $\textrm{CAG}_{i}$). The parameters ($\boldsymbol{\beta}$}, $\boldsymbol{\gamma}$}) were estimated via maximum likelihood estimation (MLE), accounting for the right-censoring in age at HD diagnosis. In addition to assuming a logistic distribution, the Langbehn model assumes that $X_{i}$ is conditionally independent of $C_{i}$ given the patient’s CAG repeats.
The Langbehn model can also provide time-to-diagnosis predictions. Building upon (1), Langbehn et al. (2004) Langbehn2004 calculated the conditional probability that patient $i$ will be diagnosed by a future $x_{f}$, given their CAG repeats $\textrm{CAG}_{i}$ and current age $x_{c}$, as
@@
assuming they have not yet been diagnosed.
Crucially, the Langbehn model does not produce a log-risk score like the other models do. However, (2) provides a single-number summary of a patient’s short-term risk of diagnosis given their age and genetic information and can be used to stratify patients by risk, which aligns the Langbehn model’s output with the other models and facilitating direct comparison. The Langbehn model cannot be estimated using existing software; we have developed R code to accompany this paper R-software.

### v2

# Accounting for Heavy Censoring in Evaluating the Risk Stratification Abilities of Existing Models for Time to Diagnosis of Huntington Disease

@@ Abstract
Huntington disease (HD) is a neurodegenerative disease with progressively worsening symptoms. Accurately modeling time to HD diagnosis is essential for clinical trial design. Langbehn’s model, the CAG-Age Product (CAP) model, the Prognostic Index Normed (PIN) model, and the Multivariate Risk Score (MRS) model have all been proposed for this task. However, these models may yield conflicting predictions and few studies have systematically compared their performance. Further, those that have could be misleading due to (i) testing the models on the same data used to train them and (ii) failing to account for high rates of right censoring ($80\%+$) in performance metrics. We discuss the theoretical foundations of these models, offering intuitive comparisons about their practical feasibility. We externally validate their risk stratification abilities using data from the ENROLL-HD study and two censoring-appropriate performance metrics, guiding the selection of a model for HD clinical trial design. As these models were originally developed in HD studies that ended more than a decade ago, we compared their predictive performance using published parameters versus updated ones (re-estimated using ENROLL-HD). We also show how these models can be used to estimate sample sizes for an HD clinical trial. Based on either metric and using published or updated parameters, the MRS model, which incorporates the most covariates, performed best. However, the much simpler PIN model offered similarly good performance while requiring fewer variables, many of which would require patients to undergo additional tests. In illustrating an HD clinical trial design, we defined an optimal threshold based on model performance metrics to determine which patients are more likely to be diagnosed. Sample size calculations using an optimal threshold based on metrics that did not account for censoring, as in previous studies, are shown to lead to underpowered trials.
Keywords Brier score $\cdot$ censored covariates $\cdot$ inverse probability of censoring weights $\cdot$ neurodegenerative disease $\cdot$ ROC curves $\cdot$ Uno’s C statistic
@@ 1.1 Huntington Disease, Genetic Testing, and Observational Studies
Huntington disease (HD) is a hereditary neurodegenerative disease caused by a repeat expansion of the trinucleotide CAG (cytosine-adenine-guanine) in the Huntingtin gene (MacDonald et al., 1993). This disease is fully penetrant for patients with $\geq 40$ CAG repeats, meaning that they will meet the diagnostic criteria for Huntington disease with probability $90\%$ (Tabrizi et al., 2022). An HD diagnosis is made based on a clinician’s assessment of a patient’s motor impairment, following the Unified Huntington Disease Rating Scale (UHDRS) (Kremer et al., 1996).
The full penetrance of HD is unique among neurodegenerative diseases. With genetic testing, researchers can identify patients who will develop HD before symptom onset and construct valuable observational cohorts (see Table 1 for examples). These studies all collect crucial clinical measurements of HD symptom severity. The UHDRS Diagnostic Confidence Level (DCL) is a categorical rating from 040-4 assigned by a clinician, where a DCL of 0 indicates no motor impairment and $4$ indicates that the clinician is $99\%$ certain that the patient is exhibiting motor impairment due to HD (Paulsen et al., 2019). Crucially, a DCL rating of $4$ indicates a clinical HD diagnosis. The UHDRS Total Motor Score (TMS) (Kremer et al., 1996) is a $31$-item, clinician-rated assessment of motor impairment, where higher scores indicate greater impairment. The Symbol Digit Modalities Test (SDMT) (Smith, 1973) and Stroop Word, Color, and Interference tests (Stroop, 1935) are timed assessments that quantify cognitive ability, with lower scores indicating worse cognitive impairment.
Study Years Sample Size Censoring Rate
COHORT 2005200520112011 $1326$ $81\%$
TRACK-HD 2008200820112011 257 $85\%$
PREDICT-HD 2002200220142014 $1159$ $77\%$
ENROLL-HD 20122012–Present $5173$ $88\%$
Table 1: Summary of major Huntington disease observational studies. COHORT (Dorsey and Investigators, 2012); TRACK-HD (Tabrizi et al., 2009); PREDICT-HD (Paulsen et al., 2006); ENROLL-HD (Sathe et al., 2021).
@@ 1.2 Designing Clinical Trials: Traditional Versus Preventative
Clinically, HD is characterized by progressive motor dysfunction, cognitive decline, and loss of functional independence. Numerous pharmacological, psychiatric, and behavioral treatments have been proposed for HD patients, but none so far have been approved for delaying or stopping disease progression. However, several studies are in progress and one has shown promise in its human trials to reduce HD progression (University College London, 2025). Another focus is on palliative care and improving quality of life, but treatments remain limited. Chorea, characterized by involuntary movements, is the only HD symptom with an approved pharmacological treatment (Olmedo-Saura, 2025). Approved symptom treatments and recent clinical breakthroughs further highlight the need for early intervention. Risk stratification models are key to designing preventative clinical trials and identifying patients before symptoms manifest, allowing for potential interventions to have the most effect before irreversible symptoms and damage are incurred.
@@
Traditionally, clinical trials for HD recruit patients who have already received their clinical diagnosis, and are thus displaying severe motor, cognitive, and functional impairment. This recruitment strategy allows for shorter trials and increases the likelihood of detecting treatment effects, which are both desirable traits to trial coordinators (Langbehn and Hersch, 2020). However, this approach has a major limitation: By the time of diagnosis (and thus trial recruitment), neurological damage is irreversible and therapeutic interventions may be ineffective. Moreover, neurodegenerative diseases like HD progress slowly for decades prior to clinical diagnosis. Thus, trials that only recruit subjects after their clinical diagnosis could fail to detect interventions that impact the disease’s early stages.
@@
Given these limitations, researchers have begun to shift clinical trial focus to patients with prodromal HD, that is, those who have not yet been diagnosed but are genetically guaranteed to develop the disease (Rubinsztein and Orr, 2016; Langbehn and Hersch, 2020). These studies, called preventative clinical trials, aim to slow or stop symptom development and thus delay or prevent a patient’s clinical HD diagnosis.
However, one unique challenge with preventative trials is the small proportion of patients expected to be diagnosed during the trial. Paulsen et al. (2019) Paulsen et al. (2019) reported that less than $1\%$ of patients in an observational study of prodromal HD patients were diagnosed within $3$ years. If recruiting patients at all HD stages based on genetic testing, we would expect similar, low incidence of diagnosis in clinical trials, as they are often just a few years in length. Any enrolled patients who remain undiagnosed at trial end will have data that were administratively censored (i.e., “cut off”) at their last follow-up visit.
@@
Since statistical power is tied to the number of “events” (HD diagnoses) in the untreated group, such trials often require large sample sizes to be sufficiently powered to detect a treatment effect. Put simply, when fewer patients are expected to be diagnosed during a trial, more must be enrolled to detect a treatment effect. With larger sample sizes, preventative trials can become more expensive and difficult to implement.
@@
Fortunately, time-to-diagnosis models can help to overcome this challenge via a strategy known as sample enrichment (Zhang et al., 2011; Long et al., 2017; Paulsen et al., 2019; Langbehn et al., 2022). We can use these models to calculate log-risk scores for potential patients and recruit only those whose log-risk scores lie above a particular threshold (i.e., who are closer to diagnosis). This approach increases the expected rate of diagnosis in the untreated group, reducing the sample size needed for a desired statistical power. Sample enrichment makes preventative clinical trials for HD more cost-efficient and easier to implement.
@@ 1.3 Comparing Existing Models for Time to Diagnosis
Time-to-diagnosis models for HD are essential tools for both clinical trial planning and patient counseling. These models help estimate when an at-risk patient (based on genetic testing) will receive a clinical diagnosis. Currently, there exist four major time-to-diagnosis models for HD: Langbehn (Langbehn et al., 2004), CAG-Age Product (CAP) (Zhang et al., 2011), Multivariate Risk Score (MRS) (Long and Paulsen, 2015), and Prognostic Index Normed (PIN) (Long et al., 2017). All four are survival models built using data from observational studies of prodromal patients ($\geq 36$ CAG repeats), as described in Section 2.2.
@@
As these models were developed, some comparisons were made to the existing models, but none considered all four. The Langbehn model was developed first, so there were no comparisons to make (Langbehn et al., 2004). The CAP model came next, and Zhang et al. (2011) Zhang et al. (2011) discussed the Langbehn model but did not provide quantitative comparisons. Long et al. (2015) Long and Paulsen (2015) compared the newly-proposed MRS model to a nonparametric model similar to CAP but again made no comparison to Langbehn. Finally, Long et al. (2017) Long et al. (2017) compared the PIN model to a semiparametric model similar to CAP, but made no comparison to Langbehn or MRS.
Despite the public availability of these modeling tools, limited work has evaluated and compared them systematically. One of the few comparisons comes from Paulsen et al. (2019) Paulsen et al. (2019), who assessed the predictive performance of the CAP, MRS, and PIN models using receiver operating characteristic (ROC) analysis. This technique evaluates how well a model separates patients who will be diagnosed up to versus after a specific time. There are three key limitations to the model evaluation conducted by Paulsen et al. (2019) Paulsen et al. (2019)
First, only three of the four major models were assessed. Second, their ROC analyses did not account for the high proportion of patients in their data who remained undiagnosed at study end (i.e., whose times to diagnosis were right censored). Although the time-to-diagnosis models were designed to handle censored data, this validation approach was not. Standard ROC analyses assume that all patients’ diagnosis times are known. Third, there was double-dipping (Ball et al., 2020), because the models were evaluated on the PREDICT-HD dataset (Paulsen et al., 2006) from which they were built. (The CAP and MRS models were developed from PREDICT-HD alone; the PIN model was developed on a combination of studies that included PREDICT-HD.)
Additional work by Long et al. (2023) Long et al. (2023) examined the PIN and CAP models in the context of the Huntington Disease Integrated Staging System (HD-ISS), which can be used as an alternative outcome measure to time to clinical diagnosis in Huntington disease clinical trials (Tabrizi et al., 2022). Long et al. (2023) Long et al. (2023) also used ROC analysis to compare model performance, incurring similar limitations to Paulsen et al. (2019) Paulsen et al. (2019) due to censoring. Moreover, their comparisons excluded the Langbehn and MRS models, and performance in predicting HD-ISS does not generalize to other HD outcomes.
@@ 2.1 Notation
Suppose we observe data for $n$ patients, indexed by i=1,,ni=1,...,n. Let $X_{i}$ denote patient $i$’s continuous diagnosis time and $\boldsymbol{Z}_{i}$ denote their covariate vector. In survival analysis, our goal is typically to estimate some aspect of the conditional distribution of $X_{i}$ given $\boldsymbol{Z}_{i}$. When the primary goal is sample enrichment, the log-risk score for $X_{i}$ given $\boldsymbol{Z}_{i}$ is the estimand of interest, denoted by ${\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}$ for some model-specific parameter vector $\boldsymbol{\beta}$} that describes this conditional distribution. This score is a scalar summary of patient $i$’s risk of being diagnosed with HD, where higher scores indicate greater risk (Park et al., 2021).
However, analysts often encounter data in which $X_{i}$ is censored by another variable $C_{i}$ for some patients. We focus on random right censoring, where $X_{i}$ is censored by the random variable $C_{i}$ if $X_{i}>C_{i}$. In many HD studies, more than $75\%$ of patients’ diagnosis times are randomly right censored (Table 1). Typically, censored patients dropped out of a study before being diagnosed or the study ended before they were diagnosed, in which case $C_{i}$ would be time to drop out or study end, respectively. Patient deaths during the trial are possible but handled in the same way as drop-outs or those censored at study end for analyses. Because preventative clinical trials focus on patients who have not yet received an HD diagnosis, any deaths observed during the trial would likely be attributed to other, non-HD causes, avoiding concerns about competing risks. Hence, we have $W_{i}=\min(X_{i},C_{i})$ and $\Delta_{i}=\textrm{I}(X_{i}\leq C_{i})$ so that our observed, patient-level data are $(W_{i},\Delta_{i},\boldsymbol{Z}_{i})$. Our goal, then, is to estimate log-risk scores ${\boldsymbol{\beta}}^{\rm T}\boldsymbol{Z}_{i}$ using these data, though our true interest is in the full data pair $(X_{i},\boldsymbol{Z}_{i})$.
@@ 2.2.1 Model 1: Langbehn
The Langbehn model was introduced to estimate the probability of HD diagnosis at a given age as a function of CAG repeats (Langbehn et al., 2004). It was developed using retrospective data collected from $40$ clinical sites across Africa, Asia, Europe, and North America. While this dataset did not come from one of the commonly cited cohorts in Table 1, it constituted the largest dataset of HD patients to date.
When developing this model, Langbehn et al. (2004) Langbehn et al. (2004) restricted the dataset to patients with $41$ to $56$ CAG repeats (inclusive). After these exclusions, they had $2913$ patients, of whom $2298$ had been clinically diagnosed by study end. With 615 patients who had not been diagnosed (i.e., whose times of diagnosis had been right-censored), the sample saw a right-censoring rate of $21\%$. This proportion contrasts sharply with the cohorts described in Table 1, which saw between $77\%$ and $88\%$ undiagnosed patients.
@@
where $\mu(\textrm{CAG}_{i};{\boldsymbol{\beta}})=\beta_{0}+\exp(\beta_{1}-\beta_{2}\textrm{CAG}_{i})$ and $\sigma^{2}(\textrm{CAG}_{i};{\boldsymbol{\gamma}})={\gamma_{0}+\exp(\gamma_{1}-\gamma_{2}\textrm{CAG}_{i})}$ are the mean age at diagnosis and its variance, respectively (both dependent on $\textrm{CAG}_{i}$). The parameters ($\boldsymbol{\beta}$}, $\boldsymbol{\gamma}$}) were estimated via maximum likelihood estimation (MLE), accounting for the right-censoring in age at HD diagnosis. In addition to assuming a logistic distribution, the Langbehn model assumes that $X_{i}$ is conditionally independent of $C_{i}$ given the patient’s CAG repeats (i.e., noninformative censoring).
The Langbehn model can also provide time-to-diagnosis predictions. Building upon (1), Langbehn et al. (2004) Langbehn et al. (2004) calculated the conditional probability that patient $i$ will be diagnosed by a future $x_{f}$, given their CAG repeats $\textrm{CAG}_{i}$ and current age $x_{c}$, as
@@
assuming they have not yet been diagnosed at age $x_{c}$.
Crucially, the Langbehn model does not produce a log-risk score like the other models do. However, (2) provides a single-number summary of a patient’s short-term risk of diagnosis given their age and genetic information and can be used to stratify patients by risk, which aligns the Langbehn model’s output with the other models and facilitating direct comparison. The Langbehn model cannot be estimated using existing software; we have developed R code to accompany this paper (R Core Team, 2022).